“Chief Grill” is a local restaurant, its annual data shows that the average sale of chicken steaks is 1200 serving per month with a standard deviation of 102. If the data is normally distributed; calculate the probability of sale of chicken steaks in the next month will be:
1) less than 1000 but not less than 1350
1
Expert's answer
2020-11-24T16:21:10-0500

Maybe more than 1000 but not less than 1350:


"P(1000<X<1350) = P(X<1350) - P(X<=1000) = F((1350-1200)\/102) - F((1000-1200)\/102) = F(1.4706) - F(-1.9608) = 0.9293 - 0.0250 = 0.9043"


Answer: 0.9043